OpenStudy (anonymous):

a wire 77 inches long is cut into two pieces whose lengths have a ratio of 5 to 2. What is the length of each piece? can you show step by step please

OpenStudy (sasogeek):

total ratio = 7

OpenStudy (sasogeek):

because 5+2=7

OpenStudy (anonymous):

that does not make any sense to me at all

OpenStudy (sasogeek):

5 out of 7 times 77=length of the longer piece, 2 out of 7 times 77=length of smaller piece

OpenStudy (sasogeek):

(5/7)*77=50inches, obviously length of smaller piece = 27 inches.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

(5/7)*77=55, not 50

OpenStudy (sasogeek):

hey mind if we talk for a bit... i'm really bored lol. u got msn or something like that :/ jw

OpenStudy (anonymous):

Making the smaller piece 22; the process is right, the last calculation is off slightly

OpenStudy (sasogeek):

i'm tired and bored, which sucks lol... hey elia.. ummm 50+22 !=77... 50+27=77 ;)

OpenStudy (anonymous):

Agreed, but (5/7)*77 is 55, not 50

OpenStudy (sasogeek):

oh my mad lol :P

OpenStudy (sasogeek):

my bad*

OpenStudy (anonymous):

5x+2x=77

OpenStudy (anonymous):

7x=77 x= 77/7 x=11

OpenStudy (sasogeek):

miscalculation... i stand to be corrected... thanks :)

OpenStudy (anonymous):

11*5=55 11*2=22 each have ,55 and 22 Medal help?

OpenStudy (sasogeek):

i gave you a medal nancy and elia

OpenStudy (anonymous):

ty

OpenStudy (sasogeek):

i'm bored :(

OpenStudy (sasogeek):

mitzirose are you okay now?

OpenStudy (anonymous):

22 and 55 am doing the minicourse also lol

OpenStudy (sasogeek):

nice

OpenStudy (sasogeek):

hey does anyone have like msn or something... i really am bored and wanna talk to someone lol... i think i've had enough math for tonight